A person with refractive myopia can see objects as close as 25 cm clearly, and objects as far away as 3 m clearly. (a) What is the maximum optical power of this person’s eye? (b) What is the minimum optical power of this person’s eye? (c) What is the range of accommodation of this person’s eye?


1
Expert's answer
2021-06-17T15:08:07-0400

Below we assume that the distance of the retina from the eye-lens is 2.5 cm.

(a) The maximum optical power of this person’s eye is


"P_1=\frac 1f=\frac{1}{v}-\frac1u=\frac{1}{0.025}-\frac{1}{-3}=40.3\text D."

(b) The minimum optical power is


"P_2=\frac{1}{v}-\frac 1u=\frac{1}{0.025}-\frac{1}{-0.25}=44\text{ D}."

(c) The range is


"\Delta P=3.7\text D"